Michael makes an investment of $100 that he knows will triple in value in 12 years. Determine the value of his investment after 30 years.
Answer by edjones(8007)

100*12x=300
1200x=300
1200x/1200=300/1200
x=1/4=.25
100*30*.25=$750. value of his investment after 30 years
.
Ed